OpenStudy (anonymous):

3. An object travelling at a speed of 40 m/s starts to accelerate uniformly with acceleration 5 m/s2. What would be the speed of the object 2 seconds later?

OpenStudy (cwrw238):

you can use one of the equations of uniform acceleration u = 40, t = 2 , a = 5, v = ? v = u + at v = 40 + 5*2

OpenStudy (anonymous):

what does u means?

OpenStudy (cwrw238):

u = initial velocity the formula can also be written as v(t) = v(0) + at v (0) = velocity at time 0 and v(t) is vel. at time t

OpenStudy (cwrw238):

- thats the better way i think the 'u' formula is whats taught in the UK

OpenStudy (anonymous):

to know the final velocity of an object after time 2 sec formula is final velocity(v)=initial velocity(u)+uniform accelration(a)*time of travelling(t) so v=u+a*t v=40+5*2 v=40+10 v=50

OpenStudy (anonymous):

1. A train is travelling with a constant acceleration of 0 m/s2. At time t = 0 s, it is travelling at a speed of 15 m/s. What is the distance travelled by the train from time t = 0 s to t = 3 s?

OpenStudy (anonymous):

for this ques how do I find the distance?

OpenStudy (anonymous):

distance(s)=initial velocity(u)*time(t)+0.5 *accelaration(a) *t *t s=15 *3+0 .5 *0 *3 *3 s=15 *3 s=45

OpenStudy (anonymous):

okay! thanks alot:)

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

its ok yar

OpenStudy (anonymous):

giv me a medal yar for my help jojoonggg